OpenStudy (anonymous):

The sum of two numbers is 66. The larger number is 2 more than 3 times the smaller. What are the two numbers? 22 and 44 20 and 46 18 and 48 16 and 50

OpenStudy (anonymous):

@Mehek14 @ganeshie8

OpenStudy (anonymous):

22 and 44?

OpenStudy (er.mohd.amir):

let numbers are a and b write equation from given data

OpenStudy (er.mohd.amir):

or try to check each option fit in given data

OpenStudy (anonymous):

idk im confused

OpenStudy (er.mohd.amir):

22*3+2=66 try check other like this

OpenStudy (er.mohd.amir):

3*a+2=b and sum a+b=66

OpenStudy (anonymous):

would the last option be my answer?

OpenStudy (anonymous):

@Mehek14 would my answer be 16 and 50?

Mehek (mehek14):

ye

OpenStudy (anonymous):

Thank u

Mehek (mehek14):

yw